Question 2: Consider the market for ice cream where the demand is given by QD 20- 2P and the supply of ice cream is given by QS 4P 10 a Graph the supply and demand curves and find the equilibrium price and quantity b Suppose the government imposes a $1 tax on ice cream, to be collected from the buyer. Plot the new curve. What is the new equilibrium price and quantity? What happens to the price paid by the buyers? What happens to the price paid by the sellers? c Who bears the greater burden of the tax? Explain why d How much tax revenue did the government generate?
Answer #1

a)

Equilibrium condition:

Demand = Supply

20 - 2P = 4P -10

30 = 6P

P = 5

Q = 20 - 2(5)

= 20 - 10

= 10

Equilibrium P = 5

Equilibrium Q = 10

Following is graph:

10 Supply Price 5 2.5 Demand 10 Quantity

b)

New Demand function after tax:

P = 10 - 0.5Q

P +1 = 10 - 0.5Q

P = 9 - .5Q

Supply:

Q = 4P - 10

P = 2.5 +0.25Q

Equilibrium:

9 - 0.5Q = 2.5 + 0.25Q

6.5 = 0.75Q

Q = 6.5/0.75

= 8.66

P = 9 -0.5(8.66)

=4.67

Old price and new quality:

P = 10 - 0.5(8.66)

=5.67

Burden on buyer: 0.67

Burden on seller: 0.33

Diagram:

10 Supply Price 5 6 Demand 2.5 8.6 10 Quantity

c)

Buyers bear the larger share of tax, since the demand by buyer is less elastic and seller equation indicates that supply is relatively elastic.

d)

Revenue collected by the government:

= Q*Tax

= 8.66*1

= $ 8.66